### v1

# Spin-Dependent Force and Inverted Harmonic Potential for Rapid Creation of Macroscopic Quantum Superpositions

Abstract
Creating macroscopic spatial superposition states is crucial for investigating matter-wave interferometry and advancing quantum sensor technology. Currently, two potential methods exist to achieve this objective. The first involves using inverted harmonic potential (IHP) to spatially delocalize quantum states through coherent inflation [1]. The second method employs a spin-dependent force to separate two massive wave packets spatially [2]. The disadvantage of the former method is the slow initial coherent inflation, while the latter is hindered by the diamagnetism of spin-embedded nanocrystals, which suppresses spatial separation. In this study, we integrate two methods: first, we use the spin-dependent force to generate initial spatial separation, and second, we use IHP to achieve coherent inflating trajectories of the wavepackets. This approach enables the attainment of massive large spatial superposition in minimal time. For instance, a spatial superposition with a mass of $10^{-15}$ kg and a size of 50 $\mu$m is realized in $0.1$ seconds. We also calculate the evolution of wave packets in both harmonic potential (HP) and IHP using path integral approach.
I Introduction
The significant interest in massive quantum superposition states primarily arises from three aspects. The first is the exploration of the quantum-classical boundary [3]. Quantum interference, from electrons to macromolecules ($10^{-31}-10^{-23}$ kg), has been observed in contemporary experiments, demonstrating their quantum nature [4, 5, 6, 7, 8]. This leads to quests about whether quantum superposition states can also be achieved for objects of larger mass [9]. The second aspect is their utility in validating theoretical models [10]. For instance, they can be employed to test wave function collapse theories [11, 12], modified quantum mechanical frameworks [13, 14, 15], and examine the weak equivalence principle [16, 17]. Additionally, they may reveal the quantum nature of gravity by combining two massive spatial superposition states [18, 19, 20, 21], see theoretical explanations  [22, 23, 24, 25, 26, 27, 28].
The third aspect is that they can act as highly sensitive quantum sensors, detecting phenomena such as the Casimir force and dipole interactions [29, 30, 31, 32], gravitational waves [33, 34], quantum sensors for detecting accelerations, and inertial rotations [35, 36, 37], dark matter [38], physics beyond the Standard Model [39], testing massive graviton [40], non-local gravitational interaction [22, 41], and analogue of light bending experiment in quantum gravity [42].
[…]
Based on these cutting-edge techniques, numerous experimental schemes for creating macroscopic spatial superposition states have been proposed [15, 2, 54, 55, 1, 56, 57, 58, 59, 60, 32]. A natural method to achieve a delocalized quantum state is to allow a quantum wave packet to evolve freely [15]. However, this delocalization process is slow. For instance, consider a silica microsphere with a mass of $10^{-15}$ kg trapped by a magnetic field at a frequency of 100 Hz and cooled to its ground state [61]. Its initial wave packet spatial width is approximately $10^{-11}$ m. After 1 s of free evolution, the wave packet width becomes about $10^{-9}$ m, roughly one-thousandth of its size. To accelerate this delocalization process, an IHP can induce coherent inflation [1, 62]. This approach allows the coherent length of a $10^{-14}$ kg nanoparticle to increase to around 1 $\mu$m in 0.6 s, making the coherence length comparable to its size. Another method for creating macroscopic spatial superposition states involves utilizing spin-dependent forces, such as using diamond embedded with a NV center [2, 54, 57, 56]. Initially, a pulse is used to place the electron spin of the NV center in a superposition state, followed by applying a magnetic field to induce spatial splitting, similar to the Stern-Gerlach experiment. The advantage of this method is the ease of preparing the initial superposition state and reading out the final spin state [63]. However, the diamagnetism of the diamond suppresses the spatial separation of the wave packets when an external magnetic field is applied [59, 64, 65, 66], limiting the direct increase of the superposition size through enhanced magnetic field gradients.
In this work, we combine spin-dependent forces and IHP to achieve a massive large spatial superposition in a relatively short time. Initially, the spin-dependent force is used to create a spatial separation between two massive wave packets. Subsequently, the IHP facilitates rapid separation of the wave packets. This approach addresses the challenges of slow acceleration in the early stages of the IHP [1, 62] and the suppression of superposition size due to diamagnetism [56, 57, 32, 59, 64, 67, 65, 66].
The paper is organized as follows. Section II presents the specific experimental protocol and the magnetic fields required to construct the HP and IHP for the experiment. Section III provides an analytical solution for the classical trajectories of the wave packets at each experiment stage. In Section IV, we numerically calculate the classical trajectories of the wave packet without approximations for the nonlinear magnetic field and compare these results with the analytical solution. Section V discusses the evolution of the wave packet under HP and IHP using path integrals. In Section VI, we examine the effect of fluctuations in the magnetic field gradient on wave packet contrast for both the HP and IHP cases. Finally, we conclude our findings in Section VII.
II Experimental scheme
The Hamiltonian of the nanodiamond embedded with a NV center in the presence of an external magnetic field is given by:
$\hat{\mathbf{H}}=\frac{1}{2m}\hat{\mathbf{P}}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}\hat{\mathbf{B}}^{2}+\hbar\gamma_{e}\hat{\mathbf{S}}\dotproduct\hat{\mathbf{B}}+\hbar D\hat{S}_{z}^{2},$ (1)
The first term represents the kinetic energy of the nanodiamond, where $\hat{\mathbf{P}}$ is the momentum and $m$ is the mass of the nanodiamond. The second term signifies the magnetic energy of a diamagnetic material (nanodiamond) in a magnetic field, with $\chi_{\rho}=-6.2\times 10^{-9}\,\,\text{m}^{3}/\text{kg}$ as the mass susceptibility and $\mu_{0}$ as the vacuum permeability. The third term is electron spin interacting with the magnetic field. $\hbar$ is the reduced Plank constant and $\gamma_{e}$ the electronic gyromagnetic ratio. $\hat{\mathbf{S}}$ is the spin operator. $\hat{\mathbf{B}}$ is the magnetic field. The last term represents the zero-field splitting of the NV center with $D=(2\pi)\times 2.8$ GHz. $\hat{S}_{z}$ is the spin component operator aligned with the NV axis. In this paper, we will assume that the nanodiamond’s rotational angular momentum is zero; see for a finite angular momentum and its implications [64]
The experimental protocol is divided into five stages: the initial separation stage, the enhancement stage, the return stage, the deceleration stage, and the recombination stage. This division aims to achieve a large spatial superposition size in a short time (around 0.1 s) and to recombine the trajectories of the CoM, thereby restoring spin coherence [51]. The five stages are outlined in Fig. 1 and described below.
Figure 1: The sketch of the experimental scheme. The light orange quadratic curves represent harmonic or IHPs. The blue and orange curves with shaded fills represent wave packets in a superposition state. The arrows inside the wave packet represent the corresponding electron spin state (spin up or spin down). The arrows along the potential indicate the direction of motion of the wave packet. The double arrows in the return stage indicate that the wave packet first separates and then returns. The time axis from top to bottom represents the initial separation stage, the enhancement stage, the return stage, the deceleration stage, and the recombination stage, or stages 1-5, respectively. $T_{1}$ to $T_{5}$ denote the evolutionary time of each stage. The grey dashed lines correspond to the moments at which the $\pi/2$ pulse is applied.
  • Initial state: The initial state is a spin superposition state (|$(\ket{+1}+\ket{-1})/\sqrt{2}$, and |$\ket{+1}$ and |$\ket{-1}$ are eigenstates with eigenvalues of 1 and -1 for the spin operator $\hat{\mathbf{S}}$.
  • Initial separation stage (Stage 1): Switching on an external linear magnetic field, the wave packets with different spin eigenstates generate spatial separation in the HP.
  • Enhancement stage (Stage 2): When the spatial separation of the two wave packets in the HP is maximized, the linear magnetic field is switched off. Then a $\pi/2$ pulse is applied to transform the state (|$(\ket{+1}+\ket{-1})/\sqrt{2}$ into |$\ket{0}$ [68, 69]. We ignore the time of applying the pulse here. Subsequently, switching on the nonlinear magnetic field creates an IHP allowing the wave packets to separate rapidly, increasing the superposition size in a short period of time.
  • Return stage (Stage 3): The nonlinear magnetic field is switched off, and the linear field is switched on after the wave packet has evolved in the IHP for a while  [^1] The exact evolution time is related to the superposition size we want to achieve. The longer the time, the larger the spatial separation between the wave packets.. The velocities of the wave packets gradually decrease in the HP and then reverse. The spatial separation between wave packets reaches its maximum at this stage.
  • Deceleration stage (Stage 4): When the magnitude of the wave packet velocity is the same as the magnitude at the end of the second stage and in the opposite direction, the linear magnetic field is switched off and the nonlinear magnetic field is switched on. The wave packet gradually decelerates in the IHP.
  • Recombination stage (Stage 5): The nonlinear magnetic field is switched off when the wave packet velocity decreases to zero. A $\pi/2$ pulse is then applied to transform the state from |$\ket{0}$ to (|$(\ket{+1}+\ket{-1})/\sqrt{2}$. The linear magnetic field is then switched on. The gradient of the magnetic field is fine-tuned so that the trajectories of the two wave packets are closed after half a period.
For clarity, we use terms like ”initial separation stage” and ”Stage 1” interchangeably without ambiguity in the following sections. To simplify the discussion, only the magnetic field component in the x-direction is considered to construct the one-dimensional HP and IHP. The HP can be obtained using a linear magnetic field [2, 56]:
$\hat{B}_{x}=B_{0}+\eta_{l}\hat{x},$ (2)
where $B_{0}$ is a bias field along the x-direction and $\eta_{l}$ is the gradient corresponding to the linear magnetic field. Substituting Eq.(2) into the Hamiltonian Eq.(1) one have:
$\hat{H}_{x}^{\text{H}}=$ $\frac{1}{2m}\hat{P}_{x}^{2}+\frac{1}{2}m\omega_{h}^{2}\hat{x}^{2}+\quantity(\hat{S}_{x}\hbar\gamma_{e}\eta_{l}-\frac{\chi_{\rho}m}{\mu_{0}}B_{0}\eta_{l})\hat{x}$
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hat{S}_{x}\hbar\gamma_{e}B_{0}+\hbar D\hat{S}_{z}^{2}.$ (3)
The superscript “H” denotes the HP. $\omega_{h}=(-\chi_{\rho}/\mu_{0})^{1/2}\eta_{l}$ is the frequency of the HP. The IHP can be obtained by means of a nonlinear magnetic field [59]:
$\hat{B}_{x}=B_{0}-\eta_{n}\hat{x}^{2},$ (4)
where $\eta_{n}$ is the gradient parameter corresponding to the nonlinear magnetic field. Note that the dimension of $\eta_{n}$ is ${\rm T/m^{2}}$. Substituting Eq.(4) into the Hamiltonian Eq.(1) one have:
$\hat{H}_{x}=$ $\frac{1}{2m}\hat{P}_{x}^{2}+\left(\frac{\chi_{\rho}m}{\mu_{0}}B_{0}\eta_{n}-\frac{\chi_{\rho}m}{2\mu_{0}}\eta_{n}^{2}\hat{x}^{2}\right)\hat{x}^{2}$
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{z}^{2}.$ (5)
Here $\hat{\mathbf{S}}\dotproduct\hat{\mathbf{B}}$ term has been ignored. This is because this term is zero during the enhancement and deceleration stages. It follows from Eq.(1) that the IHP is obtained when the condition $\expectationvalue{\hat{x}}\ll(2B_{0}/\eta_{n})^{1/2}$ is satisfied  [^2] When this condition is not satisfied, we get a quartic potential. This potential can also be used to create macroscopic spatial superposition [70]. The evolution of the wave packet in the quartic potential can be solved analytically and numerically [71, 72].. At this point, the Hamiltonian can be written as:
$\hat{H}_{x}^{\text{I}}=\frac{1}{2m}\hat{P}_{x}^{2}-\frac{1}{2}m\omega_{r}^{2}\hat{x}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{z}^{2}.$ (6)
The superscript “I” denotes the IHP. $\omega_{r}=(-2\chi_{\rho}B_{0}\eta_{n}/\mu_{0})^{1/2}$ is the frequency of the IHP.
III CoM trajectory and superposition size
The expectation value of position operator $\hat{x}$ satisfies the equation of motion:
[…]
Since the trajectories of the two wave packets are completely symmetric, to simplify the calculation process, we take the wave packet with the spin quantum number $S_{x}=1$ as an example to calculate the classical trajectory. For convenience of representation, we make the conventions shown in Table 1.
Symbol Meaning
$\omega_{i}$ The frequency of the harmonic or inverted harmonic potential at the i-th stage.
$\expectationvalue{\hat{x}}_{i}$ The classical trajectory of the i-th stage.
$X_{i}$ The classical position at the end of the i-th stage.
$\dot{X}_{i}$ The classical velocity at the end of the i-th stage.
$t_{i}$ The time variable at the i-th stage.
$T_{i}$ The time interval at the i-th stage. [^3] For example, at the beginning of the i-th stage $t_{i}=0$ and at the end of that stage $t_{i}=T_{i}$.
Table 1: The mathematical symbols that appear in calculating classical trajectories and their physical interpretations.
Stage 1 — Substituting the Hamiltonian Eq.(II) into Eq.(7) and then taking the second order derivative of the expectation value of the position operator with respect to time gives:
[…]
where $\omega_{1}$ is the frequency of the HP in the initial separation stage. The $B_{0}$ term in Eq.(8) does not affect the maximum superposition size in the initial separation stage. To be consistent with the coordinates of the later stages, we set $B_{0}$ in the initial separation stage equal to zero. Considering the initial conditions $\expectationvalue{\hat{x}(0)}=0$ and $\langle\dot{\hat{x}}(0)\rangle=0$, the solution of Eq.(8) is:
$\expectationvalue{\hat{x}}_{1}=\frac{\hbar\gamma_{e}\eta_{l}}{\omega_{1}^{2}m}(\cos(\omega_{1}t_{1})-1).$ (9)
When $t_{1}=\pi/\omega_{1}$, the superposition size achieves the maximum value $4\hbar\gamma_{e}\mu_{0}/\chi_{\rho}m\eta_{l}$ in the initial separation stage. The position of the CoM at this point is taken as the initial condition to solve the equation of motion for the enhancement stage.
Stage 2 — Using Eq.(7) again and considering the Hamiltonian in Eq.(6), one can obtain the CoM trajectory for the enhancement stage:
$\expectationvalue{\hat{x}}_{2}=X_{1}\cosh(\omega_{2}t_{2}),$ (10)
where
$X_{1}=-2\hbar\gamma_{e}\eta_{l}/\omega_{1}^{2}m.$ (11)
The $\omega_{2}$ is the frequency of the IHP in the enhancement stage.
Stage 3 — At the end of the enhancement stage, the position and velocity of the CoM are:
$X_{2}$ $=X_{1}\cosh(\omega_{2}T_{2}),$
$\dot{X}_{2}$ $=X_{1}\omega_{2}\sinh(\omega_{2}T_{2}).$ (12)
Taking the position and velocity of the CoM as the initial conditions and then combining Eq.(II) and (7) yields the trajectory of the CoM in the return stage:
$\expectationvalue{\hat{x}}_{3}$ $=X_{2}\cos(\omega_{3}t_{3})+\frac{\dot{X}_{2}}{\omega_{3}}\sin(\omega_{3}t_{3}),$
$=\sqrt{X_{2}^{2}+(\dot{X}_{2}/\omega_{3})^{2}}\sin(\omega_{3}t_{3}+\phi),$ (13)
where
$\phi=\arcsin\quantity(\frac{X_{2}}{\sqrt{X_{2}^{2}+(\dot{X}_{2}/\omega_{3})^{2}}}).$ (14)
The $\omega_{3}$ is the frequency of the HP in the return stage. The superposition size reaches its maximum value when $\sin(\omega_{3}t_{3}+\phi)=1$. The maximum superposition size is:
$\Delta X_{max}=\frac{T_{1}}{m}\frac{4\hbar\gamma_{e}}{\pi}\sqrt{\frac{\mu_{0}}{-\chi_{\rho}}}\zeta,$ (15)
where
$\zeta=\sqrt{\cosh^{2}(\omega_{2}T_{2})+\quantity(\frac{\omega_{2}}{\omega_{3}})^{2}\sinh^{2}(\omega_{2}T_{2})},$ (16)
is a dimensionless quantity. The maximum superposition size can be rewritten as:
$\Delta X_{max}\approx\quantity(\frac{3.4\times 10^{-16}\,\text{kg}}{m})\quantity(\frac{T_{1}}{\text{1 sec}})\zeta\times 10^{-6}\,\text{m}.$ (17)
The time corresponding to the maximum superposition size at the stage 3 is:
$T^{*}=\frac{1}{\omega_{3}}\quantity(\frac{\pi}{2}-\phi).$ (18)
We set the time interval of stage 3 to be $T_{3}=2T^{*}$. This is not necessary, but doing so makes the enhancement and deceleration stages symmetric. This is because, at the end of stage 3, the wave packet returns to its initial position, at which point its velocity is equal in magnitude and opposite in direction to the velocity at the beginning of the stage. We can use the same IHP as in the enhancement stage to decelerate the wave packet, thereby finally closing the wave packet trajectories. At the end of stage 3, the position and velocity of the CoM are:
$X_{3}$ $=\sqrt{X_{2}^{2}+(\dot{X}_{2}/\omega_{3})^{2}}\sin(\omega_{3}T_{3}+\phi),$
$\dot{X}_{3}$ $=\sqrt{X_{2}^{2}+(\dot{X}_{2}/\omega_{3})^{2}}\omega_{3}\cos(\omega_{3}T_{3}+\phi).$ (19)
Figure 2: The numerical calculation of trajectories. The middle plot shows the complete numerical trajectories. From left to the right they are the initial separation stage (blue), the enhancement stage (purple), the return stage (green), the deceleration stage (purple), and the recombination stage (blue). On the left and right are enlarged plots of the initial separation stage and the recombination stage, respectively. The mass $m=10^{-15}$ kg. The values of other parameters at different stages are shown in Table.2.
Stage 4 — Using the position and velocity at the end of the return stage as the initial conditions for the deceleration stage, similar to the enhancement stage, the CoM trajectory for the deceleration stage can be found by using Eq.(6) and (7) as:
$\expectationvalue{\hat{x}}_{4}=X_{3}\cosh(\omega_{4}t_{4})+\frac{\dot{X}_{3}}{\omega_{4}}\sinh(\omega_{4}t_{4}),$ (20)
where $\omega_{4}$ is the frequency of the IHP in the deceleration stage. Deriving Eq.(20) with respect to time and making it equal to zero gives the time
$t_{4}=\frac{1}{2\omega_{4}}\ln(\frac{X_{3}\omega_{4}-\dot{X}_{3}}{X_{3}\omega_{4}+\dot{X}_{3}}),$ (21)
required for the CoM velocity to decrease to zero. Substituting the evolution time $t_{4}$ into Eq.(20) gives the position of the CoM at this time
$X_{4}=\frac{1}{\omega_{4}}\sqrt{X_{3}^{2}\omega_{4}^{2}-\dot{X}_{3}^{2}}.$ (22)
If $X_{4}=0$ is assumed, then one have $\omega_{4}=-\dot{X}_{3}/X_{3}$. The reason for the negative sign is that the position is opposite in sign to the velocity at the end of return stage and $\omega_{4}$ should be greater than zero. However, Substituting this $\omega_{4}$ into $t_{4}$ gives $t_{4}\rightarrow\infty$. This is because as the CoM gets closer to the origin position, the velocity gets smaller, and at the same time the acceleration also gets smaller and eventually tends to zero. Therefore the time for the CoM to decelerate to zero tends to infinity. To avoid this situation, $X_{4}$ can only take a small value other than zero. This is why the recombination stage is needed to close the CoM trajectory.
Stage 5 — The equation of motion for the final stage (recombination stage) is the same as the initial separation stage but with a different frequency. The solution is:
$\expectationvalue{\hat{x}}_{5}$ $=-\frac{\hbar\gamma_{e}\eta_{l}}{\omega_{5}^{2}m}(\cos(\omega_{5}t_{5})+1),$
$=\frac{1}{2}X_{4}(\cos(\omega_{5}t_{5})+1),$ (23)
where $\omega_{5}$ is the frequency of the HP in the recombination stage. The second equation in Eq.(III) holds because the CoM is required to return to the origin after half a period of motion. At this point, the position and momentum of the CoM coincide.
IV Comparing analytic and numerical results
In the analytic calculation of the classical trajectories of the wave packet in Sec.III, we used approximate Hamiltonian (see Eq.(6)) in the second and fourth stages in the presence of a nonlinear magnetic field. In the numerical calculations of this section, we use the Hamiltonian without approximation (see Eq.(II)).
Stages Param. $B_{0}$(T) $\eta_{l}$(T/m) $\eta_{n}$(T/$\text{m}^{2}$) $T_{i}$(s)
1 0 2507 0.01784
2 10 $1\times 10^{6}$ 0.03000
3 0 $5\times 10^{3}$ 0.00415
4 10 $992199.56$ 0.03000
5 0 $2414.07$ 0.01853
Table 2: The values of the parameters at each stage in the calculation of the numerical trajectories. Stages 1, 3, and 5 are HPs, so $\eta_{n}$ takes no value. At these stages, $B_{0}$ takes the value 0 in order to unify the coordinate representation of each stage, but the value of $B_{0}\neq 0$, i.e. can not vanish. Stages 2 and 4 are IHPs, so $\eta_{l}$ does not take a value.
Figure 3: The superposition size and velocity differences change with time. The red solid line is the superposition size. The blue dashed line is the velocity difference. The mass $m=10^{-15}$ kg. The values of other parameters at different stages are shown in Table.2.
The numerical calculation results are shown in Fig.2. The first stage is the initial separation stage. The time of this stage is $\pi/\omega_{1}$ (half a period), which is about 0.018 s. Due to their different spin states, the two wave packets move in two different HPs and yield a spatial separation of about 6 nm. The second stage is the enhancement stage. The longer the duration of this stage, the larger the superposition size obtained. However, in order to keep the total running time around 0.1 s, we set the evolution time of this stage to be 0.03 s. With the parameters in Table.2, the spatial separation between the two wave packets at the end of the enhancement stage is about 37.14 $\mu$m. The third stage is the return stage. In this stage, the two wave packets move away from each other in the HP, and the speed of the wave packets decreases gradually. When the velocity decreases to 0, the spatial separation between them reaches a maximum value of about 50 $\mu$m. Then, the velocities of the two wave packets are reversed, and they gradually come closer together. We bring the wave packets back to roughly the initial position of this stage by fine-tuning the evolution time  [^4] The fine-tuning of the time is done here to make the trajectory more symmetric, but this tuning is not mandatory. We can also set different evolution times in this stage and then close the trajectory by adjusting the parameters of the later stages and keeping the total evolution time around 0.1 s. to 0.00415 s. The fourth stage is the deceleration stage. As analyzed in stage 4 of Sec.III, this deceleration process takes an infinite time if we want the velocity to decrease to 0 when the trajectories close. So, we make the velocity decrease to 0 when the spatial separation between the wave packets is about 6 nm by fine-tuning the magnetic field gradient. The parameter values used in this stage are shown in Table.2 and the evolution time is 0.03 s  [^5] Here, we set the evolution time to 0.03 s to be symmetric with the second stage. We can also fix the magnetic field gradient and then fine-tune the time to make the velocity of the wave packets decrease to 0 when they are separated by about 6 nm.. The fifth stage is the recombination stage. This stage is the inverse process of the initial separation stage. By fine-tuning the magnetic field gradient and evolution time in this stage, with parameters taking the values shown in Table.2, the wave packets are able to return to the initial position and with a velocity of zero.
Combining these five stages, the variations in superposition size and velocity differences are shown in Fig.3. The superposition size increases and then decreases, reaching a maximum value of 49.8294 $\mu$m at t = 0.0499 s. At this point, the velocity difference is 00, and the velocity of the wave packet starts to reverse. For comparison with the analytical expression, we substitute the parameter values from Table.2 into Eq.(15), which gives a maximum superposition state size of 49.8298 $\mu$m. The per cent error between the theoretical and numerical calculations is less than $0.001$%. This indicates that the IHP approximation we made is reasonable. By fixing the magnetic field gradient and evolution time from the first to the third stage to the values shown in Table.2, and then varying the mass, we obtain the scalar behaviour of the superposition size with respect to the mass, as shown in Fig.4.
Figure 4: The scaling behaviour of the superposition size with respect to the mass. The values of other parameters at different stages are shown in Table.2.
V Wave packet evolution
If the initial state is a Gaussian shape wave packet [^6] The difference between Gaussian wave packet and Gaussian shape wave packet is that Gaussian wave packet maintains minimum uncertainty but Gaussian shape wave packets do not necessarily maintain this property [73]. (GSWP), then the evolution of the wave packet under HP and IHP can be solved analytically [74, 75, 76, 73]. In Appendix A, we give a detailed procedure for calculating the wave packet evolution using the path integral. In this section, we provide the main results. The general form of a GSWP can be written as:
$\psi(x,t=0)=N_{0}\exp[-\frac{(x-x_{0})^{2}}{4\sigma_{0}^{2}}+i\quantity(\frac{a_{0}}{4}x^{2}+b_{0}x+c_{0})],$ (24)
where $N_{0}$ is the normalization factor, $\sigma_{0}$ is the wave packet width, $x_{0}$ is the center position of the wave packet, and $a_{0}$, $b_{0}$ and $c_{0}$ are the phase related factors. The factor 1/4 in front of the parameter $a_{0}$ is set for the convenience of later calculations. A GSWP remains a GWSP after it has evolved in HP and IHP. Their general solutions can be written as:
$\psi(x,t)=N(t)\exp[-\frac{1}{4\sigma_{x}^{2}(t)}\quantity(x-x_{c}(t))^{2}+i\quantity(\frac{a(t)}{4}x^{2}+b(t)x+c(t))].$ (25)
Where $\sigma_{x}$ is the spatial width of the wave packet evolving in time and $x_{c}(t)$ is the classical equation of motion of the wave packet. For the HP case, we have:
$\sigma_{x}^{\text{H}}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sin^{2}(\omega t)+\alpha^{2}(t))^{\frac{1}{2}},$ (26)
$x_{c}^{\text{H}}(t)$ $=\frac{\hbar a_{0}x_{0}}{2m\omega}\sin(\omega t)+x_{0}\cos(\omega t)+\frac{\hbar b_{0}}{m\omega}\sin(\omega t),$ (27)
where
$\alpha(t)=\frac{\hbar a_{0}}{2m\omega}\sin(\omega t)+\cos(\omega t)$ (28)
The index “H” indicates the expression of the physical quantity in the case of HP. For the IHP case, we have:
$\sigma_{x}^{\text{I}}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sinh^{2}(\omega t)+\beta^{2}(t))^{\frac{1}{2}},$ (29)
$x_{c}^{\text{I}}(t)$ $=\frac{\hbar a_{0}x_{0}}{2m\omega}\sinh(\omega t)+x_{0}\cosh(\omega t)+\frac{\hbar b_{0}}{m\omega}\sinh(\omega t),$ (30)
where
$\beta(t)=\frac{\hbar a_{0}}{2m\omega}\sinh(\omega t)+\cosh(\omega t)$ (31)
The index “I” indicates the expression of the physical quantity in the case of IHP. The expressions for the other parameters in Eq.(25) for the HP and IHP cases are given in Appendix A. The form of the solution of the wave packet in the IHP are the same as in the HP case, but with the replacement of ``$\sin"$ with ``$\sinh"$ and ``$\cos"$ with ``$\cosh"$. If the initial state is a Gaussian wave packet, the values of the parameters in Eq.(24) are:
$N_{0}=\frac{1}{2\pi\sigma_{0}^{2}},\,a_{0}=0,\,b_{0}=\frac{p_{0}}{\hbar}\,\,\text{and}\,\,c_{0}=-\frac{p_{0}x_{0}}{\hbar},$ (32)
where $P_{0}$ is the initial momentum. Substituting these parameters into Eqs.(26) - (31) gives:
$\sigma_{x}^{\text{H}}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sin^{2}(\omega t)+\cos^{2}(\omega t))^{\frac{1}{2}},$ (33)
$x_{c}^{\text{H}}(t)$ $=x_{0}\cos(\omega t)+\frac{p_{0}}{m\omega}\sin(\omega t),$ (34)
and
$\sigma_{x}^{\text{I}}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sinh^{2}(\omega t)+\cosh^{2}(\omega t))^{\frac{1}{2}},$ (35)
$x_{c}^{\text{I}}(t)$ $=x_{0}\cosh(\omega t)+\frac{p_{0}}{m\omega}\sinh(\omega t).$ (36)
The Eqs.(33) and (35) for the evolution of the spatial width of the wave packet in HP and IHP are the same as those in [73]. Eqs.(34) and (36) are the same as the classical equations of motion (III) and (20).
VI Fluctuation in magnetic field and wave packet contrast
Figure 5: Wave packet contrast over time when there is a small fluctuation in the linear magnetic field gradient. The red solid line denotes an initial wave packet spatial width of $\sigma_{0}=2\times 10^{-11}$ m, and the blue dashed line denotes an initial wave packet spatial width of $\sigma_{0}=3\times 10^{-11}$ m. The fluctuations in the magnetic field gradient gradually increase from left to right, for $\Delta\eta_{l}/\eta_{l}=5\times 10^{-5},1\times 10^{-4}\,\,\text{and}\,\,5\times 10^{-4}$, respectively. We set the mass $m=10^{-15}$ kg and the gradient $\eta_{l}=2507$ T/m.
In this section, we study the effect of fluctuation of the magnetic field on the contrast of the wave packets. Fluctuations in the magnetic field cause deviations from the classical position and momentum of the wave packet, preventing the two wave packets from overlapping and causing a decrease in the contrast of the wave packets. This is also known as the Humpty-Dumpty problem in the Stern-Gerlach interferometer [77].
We mark the wave packets on the two arms of the interferometer as $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$. The contrast of the wave packets can be defined as:
[…]
Since we are considering the effect of classical position and momentum deviations on the contrast, the values of the parameters are the same in $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$ except for $x_{c}(t)$ (classical position) and $b(t)$ (associated with classical momentum). The integral of Eq.(37) can be written as:
$C(t)=\exp[-\frac{\Delta x^{2}}{8\sigma_{x}^{2}(t)}-\frac{\sigma_{x}^{2}(t)\Delta b^{2}}{2}],$ (38)
where
$\Delta x$ $=x_{\text{R}}-x_{\text{L}},$
$\Delta b$ $=b_{\text{R}}-b_{\text{L}}.$ (39)
The $x_{\text{L}}$ and $x_{\text{R}}$ are the classical positions of the wave packets of the left and right arms of the interferometer. The $b_{\text{L}}$ and $b_{\text{R}}$ correspond to the values of the parameter $b$ in the left and right wave packets. For Eq.(38), only the exponential decay terms associated with the classical position and momentum deviations are retained while the amplitude and phase factor is neglected [78]. Since both the classical trajectory $x(\eta,t)$ and the parameter $b(\eta,t)$ are functions of the gradient $\eta$ [^7] For a linear magnetic field, $\eta$ is taken as $\eta_{l}$. For a nonlinear magnetic field, $\eta$ is taken as $\eta_{n}$. and time $t$, when there is a small fluctuation in the gradient, Eq.(VI) can be written as:
$\Delta x$ $=x(\eta+\Delta\eta,t)-x(\eta,t),$
$\Delta b$ $=b(p(\eta+\Delta\eta),t)-b(p(\eta),t).$ (40)
Note that the change in the parameter $b$ is caused by a change in the classical momentum, so the parameter $b$ needs to be written as an expression with respect to the momentum $p$. Substituting the expressions for $x(\eta,t)$ and $b(\eta,t)$ into Equation (VI) for the HP and IHP cases, and then combining the results with Equation (38), would reveal the effect of fluctuations in the gradient on the contrast in both cases.
VI.1 Wave packet contrast in the case of HP
The expression for the classical momentum in the HP case can be obtained by deriving Eq.(34) with respect to time:
$p(t)=-mx_{0}\omega\sin(\omega t)+p_{0}\cos(\omega t).$ (41)
Considering the Gaussian wave packet as an initial condition (see Eq.(32)) and setting the initial momentum $p_{0}=0$, in combination with Eq.(41), the parameter $b$ (see Eq.(A.1) in the Appendix A) can be written as:
$b(t)=\frac{\hbar p(t)}{4m^{2}\omega^{2}\sigma_{0}^{4}\cos^{2}(\omega t)+\hbar^{2}\sin^{2}(\omega t)}.$ (42)
Combining Eq.(34), (38), (VI), (41) and (42) gives the variation of wave packet contrast with time when there is a small fluctuation $\Delta\eta_{l}$ in the linear magnetic field gradient, as shown in Fig.5. The evolution time is set to 0.2 s to see the changing trend of the contrast more clearly. Over a short period of time, the contrast is generally oscillating and decaying with time. The larger the gradient fluctuation, the faster the decay. The initial wave packet spatial width also has an effect on the contrast. The wider the initial spatial width, the slower the contrast decays.
VI.2 Wave packet contrast in the case of IHP
Figure 6: Wave packet contrast over time when there is a small fluctuation in the nonlinear magnetic field gradient. The red, blue, and orange solid (dashed) lines represent gradient fluctuations of $\Delta\eta_{n}/\eta_{n}=1\times 10^{-4},3\times 10^{-4}\,\,\text{and}\,\,5\times 10^{-4}$, respectively. The solid lines correspond to an initial wave packet spatial width of $\sigma_{0}=2\times 10^{-11}$ m, and the dashed lines correspond to an initial wave packet spatial width of $\sigma_{0}=3\times 10^{-11}$ m. We set the mass $m=10^{-15}$ kg and the gradient $\eta_{n}=1\times 10^{6}$ T/$\text{m}^{2}$.
Derivation of Eq.(36) with respect to time yields the momentum for the IHP case:
$p^{\prime}(t)=mx_{0}\omega\sinh(\omega t)+p_{0}\cosh(\omega t).$ (43)
Consider also the Gaussian wave packet as an initial condition (see Eq.(32)) and set the initial momentum $p_{0}=0$. Combined with Eq.(43), the parameter $b$ (see Eq.(A.2) in the Appendix A) can be written as:
$b(t)=-\frac{\hbar p^{\prime}(t)}{4m^{2}\omega^{2}\sigma_{0}^{4}\cosh^{2}(\omega t)+\hbar^{2}\sinh^{2}(\omega t)}.$ (44)
Combining Eq.(36), (38), (VI), (43) and (44) gives the variation of wave packet contrast with time when there is a small fluctuation $\Delta\eta_{n}$ in the nonlinear magnetic field gradient, as shown in Fig.6. The wave packet contrast decays with time. The smaller the fluctuation $\Delta\eta_{n}$, the slower the decay. Also, increasing the initial wave packet spatial width can slow down the wave packet contrast decay.
For both HP and IHP cases, reducing the fluctuation of the magnetic field gradient and increasing the initial wave packet width can suppress the decay of contrast.
VII Conclusion
In this paper, we investigated the use of spin-dependent force and IHP to generate massive large spatial superposition states and construct a full-loop interferometer. The spin-dependent force enables a massive wave packet to achieve a small spatial separation in a short time. Subsequently, the IHP allows this separation to increase rapidly. For a nanodiamond with a mass of $10^{-15}$ kg, we found that a superposition size of 50 $\mu$m can be achieved in 0.1 s. We also analytically calculated the evolution of the wave packet in both the HP and IHP using path integrals. Additionally, we discussed how fluctuations in the magnetic field gradient affect the wave packet contrast, concluding that smaller gradient fluctuations and a wider wave packet width result in a slower decrease in contrast.
Our study focused on a nanodiamond embedded with an NV centre. The nanodiamond possesses mechanical degrees of freedom (CoM motion), internal degrees of freedom (phonons), and rotational degrees of freedom. Currently, both center-of-mass motion and phonons can be cooled to the ground state [49], and phonons are difficult to excite during movement in HP and IHP [79, 80]. The rotational degree of freedom can affect the final wave packet contrast of the interferometer; however, if the NV center is at or near the nanodiamond center, this effect can be mitigated by fine-tuning the magnetic field and timing [81, 82, 64]. Furthermore, the electron spin coherence time of the NV center can be as long as 1 s at low temperatures and under pure nanocrystal conditions [83, 84, 85]. These findings indicate the feasibility of realizing macroscopic quantum states with masses ranging from $10^{-17}$ to $10^{-14}$ kg and superposition sizes on the order of micrometres in the laboratory.
Note added: Following the completion of our work, we identified corresponding independent research by Sougato Bose’s group (on arXiv today).
acknowledgments
We are grateful to Tian Zhou, Ryan Marshman and Sougato Bose for stimulating discussions. R. Z. and Q. X. are supported by China Scholarship Council (CSC) fellowship. A. M. is supported by Sloan and the Gordon and Betty Moore foundations.
Appendix A Wave packet evolution
According to the path integral, the evolution of the wave function can be written as:
[…]
where $\psi(x^{\prime},0)$ represents the wave function at the initial moment. $K(x,t;x^{\prime},0)$ is the propagator. When the potential energy is quadratic, the propagator can be calculated by the Van Vleck-Pauli-Morette formula:
[…]
where $S_{c}$ is the classical action quantity and defined as:
[…]
$\mathcal{L}$(t) is the Lagrangian of the system. Assuming that the solution to the classical trajectory of the system is $x_{c}(t)$, the Lagrangian can be written as:
$\mathcal{L}(t)=\frac{1}{2}m\dot{x}^{2}(t)-\frac{1}{2}m\omega^{2}x^{2}(t).$ (48)
A.1 Wave packet evolution in a HP
The general solution for the classical trajectory of a wave packet in a HP is:
$x(t)=x_{0}\cos(\omega t)+\frac{p_{0}}{m\omega}\sin(\omega t),$ (49)
where $x_{0}$ and $p_{0}$ are the classical initial position and initial momentum of the wave packet, respectively. Consider the boundary conditions $x(t_{i}=0)=x_{i}$ and $x(t_{f})=x_{f}$. Substituting them into Eq.(49) yields:
$x(t)=x_{i}\cos(\omega t)+\frac{x_{f}-x_{i}\cos(\omega t_{f})}{\sin(\omega t_{f})}\sin(\omega t).$ (50)
Combining Eq.(47), (48) and (50) gives the classical action at the harmonic potential as:
$S=\frac{m\omega}{2}\frac{(x_{f}^{2}+x_{i}^{2})\cos(\omega t)-2x_{f}x_{i}}{\sin(\omega t)}.$ (51)
Note that after integrating in Eq.(47), the time parameter in the action is “$t_{f}$”. In Eq.(51) we replace “$t_{f}$” with “$t$”, thus aligning with the time variable in Eq.(45). Substituting Eq.(51) into Eq(46) results in the propagator of the wave packet at the HP as:
$K(x_{f},t;x_{i},0)=\sqrt{\frac{m\omega}{i2\pi\hbar\sin(\omega t)}}\exp[\frac{i}{\hbar}\frac{m\omega}{2}\frac{(x_{f}^{2}+x_{i}^{2})\cos(\omega t)-2x_{f}x_{i}}{\sin(\omega t)}].$ (52)
Since both the initial wave function (Eq.(24)) and the propagator (Eq.(52)) are Gaussian quadratic functions, solving Eq.(45) for the wave packet evolution is a Gaussian quadratic integral. The result of the integration is still a Gaussian quadratic function:
$\psi(x,t)=N(t)\exp[i\frac{x^{2}}{4u_{t}^{2}}-\frac{x_{0}^{2}}{4\sigma_{0}^{2}}]\exp[\frac{\quantity(ib_{0}-i\frac{x}{2u_{t}^{2}\cos(\omega t)}+\frac{x_{0}}{2\sigma_{0}^{2}})^{2}}{\frac{1}{\sigma_{0}^{2}}-i\quantity(\frac{1}{u_{t}^{2}}+a_{0})}],$ (53)
where
$u_{t}^{2}$ $=\frac{\hbar\sin(\omega t)}{2m\omega\cos(\omega t)},$
$N(t)$ $=N_{0}\sqrt{\frac{m\omega}{i2\pi\hbar\sin(\omega t)}}\sqrt{\frac{4\pi}{1/\sigma_{0}^{2}-i\quantity(1/u_{t}^{2}+a_{0})}}e^{ic_{0}}.$ (54)
Eq.(53) can be rewritten in the familiar form of the GSWP:
$\psi(x,t)=N(t)\exp[-\frac{1}{4\sigma_{x}^{2}(t)}\quantity(x-x_{c}(t))^{2}+i\quantity(\frac{a(t)}{4}x^{2}+b(t)x+c(t))],$ (55)
where
$\sigma_{x}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sin^{2}(\omega t)+\quantity(\frac{\hbar a_{0}}{2m\omega}\sin(\omega t)+\cos(\omega t))^{2})^{\frac{1}{2}},$ (56)
$x_{c}(t)$ $=\frac{\hbar a_{0}x_{0}}{2m\omega}\sin(\omega t)+x_{0}\cos(\omega t)+\frac{\hbar b_{0}}{m\omega}\sin(\omega t),$ (57)
which represent the spatial width of the wave packet and the classical equation of motion of the nanoparticle, respectively. The expressions for the three parameters in the imaginary part are:
$a(t)$ $=\frac{1}{u_{t}^{2}}-\frac{1+a_{0}u_{t}^{2}}{4u_{t}^{6}\cos^{2}(\omega t)\quantity(\quantity(1/u_{t}^{2}+a_{0})^{2}+1/\sigma_{0}^{4})},$
$b(t)$ $=\frac{2b_{0}\sigma_{0}^{4}-u_{t}^{2}\quantity(x_{0}-2a_{0}b_{0}\sigma_{0}^{4})}{2\cos(\omega t)\quantity(\sigma_{0}^{4}+2a_{0}u_{t}^{2}\sigma_{0}^{4}+u_{t}^{4}\quantity(1+a_{0}^{2}\sigma_{0}^{4}))},$
$c(t)$ $=\frac{x_{0}^{2}+x_{0}u_{t}^{2}(4b_{0}+a_{0}x_{0})-4b_{0}^{2}\sigma_{0}^{4}(1+a_{0}u_{t}^{2})}{4u_{t}^{2}\sigma_{0}^{4}\quantity(\quantity(1/u_{t}^{2}+a_{0})^{2}+1/\sigma_{0}^{4})}.$ (58)
A.2 Wave packet evolution in a IHP
The calculation process for the evolution of the wave packet in the IHP is the same as in the case of the HP. The form of the classical equation of motion and the action of the wave packet in the IHP are the same as in the HP case, but with the replacement of ``$\sin"$ with ``$\sinh"$ and ``$\cos"$ with ``$\cosh"$. According to Eq.(46), the propagator at the IHP is obtained as:
$K^{\prime}(x_{f},t;x_{i},0)=\sqrt{\frac{m\omega}{i2\pi\hbar\sinh(\omega t)}}\exp[\frac{i}{\hbar}\frac{m\omega}{2}\frac{(x_{f}^{2}+x_{i}^{2})\cosh(\omega t)-2x_{f}x_{i}}{\sinh(\omega t)}].$ (59)
Using Eq.(45) again, multiplying this propagator with the initial wave function and integrating over the initial position gives:
$\psi^{\prime}(x,t)=N^{\prime}(t)\exp[i\frac{x^{2}}{4v_{t}^{2}}-\frac{x_{0}^{2}}{4\sigma_{0}^{2}}]\exp[\frac{\quantity(ib_{0}-i\frac{x}{2v_{t}^{2}\cosh(\omega t)}+\frac{x_{0}}{2\sigma_{0}^{2}})^{2}}{\frac{1}{\sigma_{0}^{2}}-i\quantity(\frac{1}{v_{t}^{2}}+a_{0})}],$ (60)
where
$v_{t}^{2}$ $=\frac{\hbar\sinh(\omega t)}{2m\omega\cosh(\omega t)},$
$N^{\prime}(t)$ $=N_{0}\sqrt{\frac{m\omega}{i2\pi\hbar\sinh(\omega t)}}\sqrt{\frac{4\pi}{1/\sigma_{0}^{2}-i\quantity(1/v_{t}^{2}+a_{0})}}e^{ic_{0}}.$ (61)
Rearranging Eq.(60) yields:
$\psi^{\prime}(x,t)=N^{\prime}(t)\exp[-\frac{1}{4\sigma_{x}^{\prime 2}(t)}\quantity(x-x^{\prime}_{c}(t))^{2}+i\quantity(\frac{a^{\prime}(t)}{4}x^{2}+b^{\prime}(t)x+c^{\prime}(t))],$ (62)
where
$\sigma_{x}^{\prime}(t)$ $=\sigma_{0}\quantity(\frac{\hbar^{2}}{4m^{2}\omega^{2}\sigma_{0}^{4}}\sinh^{2}(\omega t)+\quantity(\frac{\hbar a_{0}}{2m\omega}\sinh(\omega t)+\cosh(\omega t))^{2})^{\frac{1}{2}},$ (63)
$x^{\prime}_{c}(t)$ $=\frac{\hbar a_{0}x_{0}}{2m\omega}\sinh(\omega t)+x_{0}\cosh(\omega t)+\frac{\hbar b_{0}}{m\omega}\sinh(\omega t),$ (64)
which are the spatial width of the wave packet and the classical equation of motion at the IHP. The expressions for the parameters of the imaginary part are:
$a^{\prime}(t)$ $=\frac{1}{v_{t}^{2}}-\frac{1+a_{0}v_{t}^{2}}{4v_{t}^{6}\cosh^{2}(\omega t)\quantity(\quantity(1/v_{t}^{2}+a_{0})^{2}+1/\sigma_{0}^{4})},$
$b^{\prime}(t)$ $=\frac{2b_{0}\sigma_{0}^{4}-v_{t}^{2}\quantity(x_{0}-2a_{0}b_{0}\sigma_{0}^{4})}{2\cosh(\omega t)\quantity(\sigma_{0}^{4}+2a_{0}v_{t}^{2}\sigma_{0}^{4}+v_{t}^{4}\quantity(1+a_{0}^{2}\sigma_{0}^{4}))},$
$c^{\prime}(t)$ $=\frac{x_{0}^{2}+x_{0}v_{t}^{2}(4b_{0}+a_{0}x_{0})-4b_{0}^{2}\sigma_{0}^{4}(1+a_{0}v_{t}^{2})}{4v_{t}^{2}\sigma_{0}^{4}\quantity(\quantity(1/v_{t}^{2}+a_{0})^{2}+1/\sigma_{0}^{4})}.$ (65)